OpenStudy (anonymous):

( cot x)' |x= pi/4 solve

OpenStudy (anonymous):

What is stopping you?

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

Put yourself in my situation. If someone told you they don't know how to sit down, you'd ask them "why not", right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok so i got the diffrnciation of cot x , its csc^2 (x) but what shoud i do with pi/4?

OpenStudy (anonymous):

Substitute x with pi/4

OpenStudy (anonymous):

so what will we get ? 2 ?

OpenStudy (anonymous):

hold on let me think...

OpenStudy (anonymous):

First of all, shouldn't it be \((\cot x) '=-\csc ^2 x\)

OpenStudy (anonymous):

Second of all, \(\csc x=1/(\sin x)\). \(\sin(\pi/4) = \sqrt 2 / 2\)

OpenStudy (anonymous):

wolfram alfa says 2

OpenStudy (anonymous):

Hmmm, are you trying to irritate me?

OpenStudy (anonymous):

no lol

OpenStudy (anonymous):

Is the purpose of this question to look things up on wolfram alpha, or is it get the answer using math?

OpenStudy (anonymous):

I haven't even given you an answer, only facts that would help you find an answer.

OpenStudy (anonymous):

i know , can u give me answer

OpenStudy (anonymous):

Well, I'd like to have some evidence that you attempted the problem first.

OpenStudy (anonymous):

If so, I wouldn't mind giving you the answer and how to get the answer.

OpenStudy (anonymous):

ok let me do it

OpenStudy (anonymous):

ok so whats next ?

OpenStudy (anonymous):

u there ?